OpenStudy (anonymous):

find the exact value of cot(11pi/6) cos(11pi/6) csc(11pi/6)

OpenStudy (anonymous):

can I have the full solution please? thankyou!

hartnn (hartnn):

first remember that 11pi/6 is in 4th quadrant, so only cos and sec ratio will be positive... cot(11pi/6)=-cot (2pi-11pi/6)---->as cot x=-cot(-x)=-cot (2pi-x) =-cot(pi/6)=- sqrt(3) now cos(11pi/6)=cos (2pi-11pi/6)---->as cos x=cos(-x)=cos (2pi-x) =cos(pi/6)=sqrt(3)/2 csc(11pi/6)=-csc (2pi-11pi/6)---->as csc x=-csc(-x)=-csc (2pi-x) =-csc(pi/6)=- 2

hartnn (hartnn):

@ienarancillo

OpenStudy (anonymous):

 is this correct? thanks